A blender requires 350 W to power the rotating blades that chop food. If the blender has a resistance of 75, how much current passes through the blender’s circuit? What is the potential difference across the blender?
1
Expert's answer
2020-06-08T10:27:44-0400

We know that the power of the blender can be expressed in terms of current and resistance as


"P=I^2R\rightarrow I=\sqrt{\frac{P}{R}}=2.2\text{ A}."

The potential difference, or voltage, can be calculated by Ohm's law:


"V=IR=162\text{ V}."